# System performance of a cryogenic test-bed for the time-division multiplexing readout for NewAthena X-IFU

### v2

@@ 3.3 Bandwidth optimization
As mentioned in Section 2.1, the old flex design resulted in a large parasitic inductance which caused underdamped oscillations (“ringing”) in the electrical signals of the systems. Such ringing is expected to be initiated by two mechanisms: the switching transients sequentially activating the TDM rows and large change in the feedback signal, such as when the readout switches to a row with an X-ray event. This mechanism can be characterized by injecting a signal to one SQ1 feedback coil of amplitude comparable to the height of a 6 keV event, which for this system is at the level of 2, with Wb being the magnetic flux quantum and representing a single period in the vs. characteristic curve. The observed response for the old flex is plotted in black in Figure 5. We measured a peak frequency is such oscillation at a frequency of MHz, well within the system bandwidth. In TDM systems with regular short harness, such ringing is present but at a higher frequency that is outside the typical 12-15 MHz readout bandwidth, allowing for 160 ns row time readout.
With the old flex harness, the system performance could be retained by operating with an increased row time of 360 ns, corresponding to a line rate of 2.7 MHz and sampling for 64 ns instead of 32 ns. As a consequence, to maintain roughly the same frame time, the multiplexing factor was reduced from 32 to 16. With this configuration, both at NIST and at SRON, similar systems were able to achieve average energy resolution of 2.8 eV FWHM [14].
Subsequently, the new flex design was implemented in the system. A ringing characterization measurement was repeated, by injecting a 2 signal at the SQ1 feedback coil. In this case (Figure 5, green curve), no underdamped oscillations were observed. However, most likely due to the larger capacitance to ground in the new design, a significantly reduced bandwidth was observed, at the level of 5 MHz.
In consequence of this, the decay time of the transient was at a level of 200 ns, longer than the 160 ns row time, which made the use of the nominal timing settings infeasible. This effect on the switching transient due to finite analog bandwidth is known and consistent with what is reported in Durkin et al. (2021) [21], and it is equivalent to applying an aliasing filter to the FB1 signal path.
We further verified this by simulating the system behavior in a Dymola model. A low pass filter was inserted in the feedback path to simulate the limited bandwidth of the flexes and the cut-off was varied to several values. The impact on the transients is reported Figure 6, from which a cut-off at a level of 5 MHz is qualitatively consistent with the experimental observations.
@@ 4 System performance demonstration
To measure the X-ray performance of the system, about 10’000 X-ray events per pixel are collected from the Fe-55 source installed on the snout’s chimney, with a count rate of approximately 1.2 counts per second per pixel. The energy of the photons is estimated using the optimal filtering technique[23]. To assess the spectral performance, the MnK model [24] is fitted to the collected, calibrated data. The model is convolved with a Gaussian broadening to parametrize the FWHM energy resolution of the detectors. The goodness of the fit is estimated with the maximum likelihood method by means of a statistics modified to mitigate biases due to bin size or number of events[25].
Each channel is activated by biasing the SSA, SQ1 and TES detectors on their optimal operating point. For the SSA, this is done by measuring its V- curves for different bias currents and selecting the one with the largest modulation depth. The flux offset is chosen to operate the SSA on the steeper slope of its vs. curve, both to minimize the system noise and to maximize the readout bandwidth. The same is repeated for the bias current of the SQ1s, selecting the one producing the maximum modulation depth. The flux offset of the SSA is then varied iteratively per each row to find the value maximising each SQ1s modulation depth. A voltage offset is then applied to center the FLL signal around zero, to optimize the feedback DAC dynamic range.
The TES bias is set to a value corresponding to roughly 17% of its normal resistance value, has shown to yield on average the best spectral performance, at the 150 nT magnetic field offset.
@@
Time-neighbor cuts tag events which happen within a certain time window from an event happening in a subsequent or preceding row, while global cuts tag events happening within a time window regardless of their row number. In both cases, the cuts are limited to within the same Column. To simplify the data analysis pipeline, we have applied here a crosstalk-rejection strategy that cuts all perpetrator-victim pairs within a readout Column, with a time window of 2 ms. This strategy is suboptimal with respect to what is being envisaged for X-IFU, so that the throughput of surviving counts in this data set should not be understood as an analysis of what is possible for the flight pipeline of X-IFU. Furthermore, the impact of crosstalk on in-flight performance is expected to be lower than what measured in the laboratory with the approximately monochromatic Fe-55 source, since the typical models assume a Crab-like energy spectrum, with an average energy lower than 6 keV [11, 27].
Experiments with 1.2 cps count rate showed a larger impact on the co-added system performance. However, typical TDM laboratory demonstrations use a count rate at a level of 0.3 cps per pixel, similar to the expected average count rate in for X-IFU when not observing high intensity sources such as supernova remnants, gamma ray burst afterglows or XRBs (X-ray binaries) in our galaxy. We therefore repeated the experiment by reducing the count rate to about 0.7 cps per pixel, by increasing the thickness of the Al foil in front of the cold 55Fe source.
@@ 5 Conclusions and future outlook
This demonstration represents a fundamental progress for the FPA-DM1.1 of X-IFU as it is the first time that a TDM system is operated with 1.5 m long harness, closer to the X-IFU’s 2.3 m requirements.
NewAthena’s instrument will employ twisted-pair looms in balanced differential operation, and so will the FPA-DM1.1. Having reached the expected level of performance for the system, in the near future we will replace the current flex harness with twisted-pair looms, to assess their impact not only on performance but also on crosstalk and analog bandwidth. This will be an important intermediate step before the integration and test campaign of the X-IFU FPA-DM1.1, foreseen for Q1 2025. The comparison of the system behavior with flex harness and twisted-pair looms will be reported on a separate paper.

### v3

@@ 3.3 Bandwidth optimization
As mentioned in Section 2.1, the old flex design resulted in a large parasitic inductance which caused underdamped oscillations (“ringing”) in the electrical signals of the systems. Such ringing is expected to be initiated by two mechanisms: the switching transients sequentially activating the TDM rows and large changes in the feedback signal, such as when the readout switches to/from a row reading out a TES hit by an X-ray.
This mechanism can be characterized by injecting to the feedback coil of one SQ1 (first-stage SQUID) a signal of amplitude comparable to the one caused by an X-ray releasing energy in the TES. For this system, the pulse height for a 6 keV photon roughly corresponds to a signal of 2, with Wb being the magnetic flux quantum and representing a single period in the vs. characteristic curve.
This experiment was performed first for the old flex and the observed response is plotted in blue in Figure 5. We measured a peak amplitude is such oscillation at a frequency of MHz, well within the system bandwidth. In TDM systems with a regular short harness, such ringing is present but at a higher frequency that is outside the typical 12-15 MHz readout bandwidth, allowing for 160 ns row time readout.
With the old flex harness, the system performance could be retained by operating with an increased row time of 360 ns, corresponding to a line rate of 2.7 MHz and sampling for 64 ns instead of 32 ns. As a consequence, to maintain roughly the same frame time, the multiplexing factor was reduced from 32 to 16. With this configuration, both at NIST and at SRON, similar systems were able to achieve an average energy resolution of 2.8 eV FWHM [15].
Subsequently, the new flex design was implemented in the system. A ringing characterization measurement was repeated, by injecting a 2 signal at the SQ1 feedback coil. In this case (Figure 5, dashed orange curve), no underdamped oscillations were observed. However, most likely due to the larger capacitance to ground in the new design, a significantly reduced bandwidth was observed, at the level of 5 MHz.
In consequence of this, the decay time of the transient was at a level of 200 ns, longer than the 160 ns row time, which made the use of the nominal timing settings infeasible. This effect on the switching transient due to finite analog bandwidth is known and consistent with what is reported in Durkin et al. (2021) [22], and it is equivalent to applying an aliasing filter to the SQ1 feedback signal path.
We further verified this by simulating the system behavior in a Dymola[24] model. A low pass filter was inserted in the feedback path to simulate the limited bandwidth of the flexes and the cut-off was varied to several values. The impact on the transients is reported Figure 6, from which a cut-off at a level of 5 MHz is qualitatively consistent with the experimental observations.
@@ 4 System performance demonstration
To measure the X-ray performance of the system, about 10,000 X-ray events per pixel are collected from the Fe-55 source installed on the snout’s chimney, with a count rate of approximately 1.2 counts per second per pixel. The energy of the photons is estimated using the optimal filtering technique[25]. To assess the spectral performance, the MnK model [26] is fitted to the collected, calibrated data. The model is convolved with a Gaussian broadening to parametrize the FWHM energy resolution of the detectors. The goodness of the fit is estimated with the maximum likelihood method by means of a statistics modified to mitigate biases due to bin size or number of events[27].
Each channel is activated by biasing the SSA, SQ1 and TES detectors on their optimal operating point. For the SSA, this is done by measuring its V- curves for different bias currents and selecting the one with the largest modulation depth. The flux offset is chosen to operate the SSA on the steeper slope of its vs. curve, both to minimize the system noise and to maximize the readout bandwidth. The same is repeated for the bias current of the SQ1s, selecting the one producing the maximum modulation depth. The flux offset of the SSA is then varied iteratively per each row to find the value maximising each SQ1’s modulation depth. A voltage offset is then applied to center the FLL signal around zero, to make optimal use of the feedback DAC dynamic range.
The TES bias is set to a value corresponding to roughly 17% of its normal resistance value, which during early measurements had shown to yield on average the best spectral performance, at the 150 nT magnetic field offset.
@@
Time-neighbor cuts tag events which happen within a certain time window from an event happening in a subsequent or preceding row, while global cuts tag events happening within a time window regardless of their row number. In both cases, the cuts are limited to within the same Column. To simplify the data analysis pipeline, we have applied here a crosstalk-rejection strategy that cuts all perpetrator-victim pairs within a readout Column, with a time window of 2 ms. This strategy is suboptimal with respect to what is being envisaged for X-IFU, so that the throughput of surviving counts in this data set should not be understood as an analysis of what is possible for the flight pipeline of X-IFU. Furthermore, the impact of crosstalk on in-flight performance is expected to be lower than what is measured in the laboratory with the approximately monochromatic Fe-55 source, since the typical models assume a Crab-like energy spectrum, with an average energy lower than 6 keV [11, 29].
Experiments with a 1.2 cps count rate showed a larger impact on the co-added system performance. However, typical TDM laboratory demonstrations use a count rate at a level of 0.3 cps per pixel, similar to the expected average count rate for X-IFU when not observing high intensity sources such as supernova remnants, gamma ray burst afterglows or XRBs (X-ray binaries) in our galaxy. We therefore repeated the experiment by reducing the count rate to about 0.7 cps per pixel, by increasing the thickness of the Al foil in front of the cold 55Fe source.
@@ 5 Conclusions and future outlook
This demonstration represents a fundamental progress for the FPA-DM1.1 of X-IFU as it is the first time that a TDM system is operated with a 1.5 m long harness, closer to the X-IFU’s 2.3 m requirements.
NewAthena’s instrument will employ twisted-pair looms in balanced differential operation, and so will the FPA-DM1.1. Having reached the expected level of performance for the system, in the near future we will replace the current flex harness with twisted-pair looms, to assess their impact not only on performance but also on crosstalk and analog bandwidth. This will be an important intermediate step before the integration and test campaign of the X-IFU FPA-DM1.1, foreseen for Q2 2025, which will allow to assess the TDM readout operation in a configuration more representative of X-IFU. The comparison of the system behavior with a flex harness versus twisted-pair looms will be reported on a separate paper.
@@ Acknowledgements
Conflict of interest
The authors declare there are no financial interests, commercial affiliations, or other potential conflicts of interest that have influenced the objectivity of this research or the writing of this paper.